OpenStudy (anonymous):

how can I solve: 1+1/2 all over 2+1/4 im not sure when I have to find the common denominator, before of after i flip the fractions since it's devision. or do i not flip?

OpenStudy (anonymous):

so: \[\frac{ 1+\frac{ 1 }{ 2 } }{ 2+\frac{ 1 }{ 4 } }\] well convert those mixed numbers to improper fractions: \[\frac{ \frac{ 3 }{ 2 } }{ \frac{ 9 }{ 4 } }\] that equals; \[\frac{ 3 }{ 72 }\] because a/b/c/d = a/(b*c*d) so i multiplied 2*9*4 which is 72 so you get 3/72 if you simplify that by dividing by 3 on top and bottom you get 1/24 ANSWER: 1/24

OpenStudy (mathmale):

If I read your question correctly, you want to EVALUATE (1+1/2)/ (2 + 1/4). Is that correct? If so, why not convert 1 + 1/2 = 1 1/2 to a proper fraction, meaning 3/2? And then do the same thing to the denominator. Finally, flip over the fraction from the denominator and multiply the numerator by this result. mihirb has done a very good job of explaining the solution to this problem.

OpenStudy (anonymous):

the answer says it's 2/3 however

OpenStudy (anonymous):

how can i solve that by finding a common denominator? we're doing rational proportions in class